At what distance above the surface of the earth is the acceleration due to the earth’s gravity 0.980 π‘š/𝑠 2 if the acceleration due to gravity at the surface has a magnitude of 9.80 π‘š/𝑠 2 ?


1
Expert's answer
2021-12-20T10:26:31-0500

"g=F/m=GM/r^2\to r=\sqrt{GM/g}="


"=\sqrt{6.67\cdot 10^{-11}\cdot 5.97\cdot10^{24}/0.98}="


"=20157518\ (m)"


"h=20157518-6371000=13786518\ (m)"